Martina leaves home at 9 a.m., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
Answer by stanbon(75887)

Martina leaves home at 9 a.m., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
-----------
Martina DATA:
Rate = 24 mph ; Distance = x miles ; Time = d/r = x/24 hrs
------------------
John DATA:
Rate = 48 mph ; Distance = x miles ; Time = d/r = x/48 hrs
-------------------
EQUATION:
Martina time = John time - 2 hrs
x/24 = x/48 -2
Multiply thru by 48 to get:
2x = x - 96
x = 96 miles
----------------
Martina time = 96/24 = 4 hrs
9:00 + 4 hrs = 1:00 PM
------------------
John will catch up to Martica at 1:00 PM
================
Cheers,
Stan H.